how many birds and lions do you have in your zoo? If there are 30 heads and 100 feet.
Answer by stanbon(48546)

how many birds and lions do you have in your zoo? If there are 30 heads and 100 feet.
----------
Let # of birds be "b"; let number of lions be "l"
Head equation: b+l = 30
Feet equation: 2b+4l = 100
--------------
l = 30-b
Substitute to get:
2b + 4(30-b) = 100
Solve for b:
2b+120-4b = 100
-2b = -20
b = 10 (# of birds)
l = 30 - 10 = 20 ( # of lions)
===============
Cheers,
Stan H.